Please help me with this problem: Lee worked 22 hours this week and made $132. If she works 15 hours next week at the same pay rate, how much will she make?
Found 2 solutions by MRperkins, stanbon:
Answer by MRperkins(300)

Join me in an online whiteboard to see this problem worked out. Email me at justin.sheppard.tech@hotmail.com and I will send you the link for the online whiteboard.
Respectfully,
Justin

Answer by stanbon(75887)

Lee worked 22 hours this week and made $132. If she works 15 hours next week at the same pay rate, how much will she make?
------------
Proportion:
x/(15 hr) = $132/(22 hr)
x = 15(132/22)
x = 15*66/11
x = $90.00
==================
Cheers,
Stan H.